# Accelerating Monte–Carlo Tree Search with Optimized Posterior Policies

### v1

@@ Abstract.
We introduce a recursive AlphaZero-style Monte–Carlo tree search algorithm, “RMCTS”. The advantage of RMCTS over AlphaZero’s MCTS-UCB [silver2018general] is speed. In RMCTS, the search tree is explored in a breadth-first manner, so that network inferences naturally occur in large batches. This significantly reduces the GPU latency cost. We find that RMCTS is often more than 40 times faster than MCTS-UCB when searching a single root state, and about 3 times faster when searching a large batch of root states.
The recursion in RMCTS is based on computing optimized posterior policies at each game state in the search tree, starting from the leaves and working back up to the root. Here we use the posterior policy explored in “Monte–Carlo tree search as regularized policy optimization” [grill2020]. Their posterior policy is the unique policy which maximizes the expected reward given estimated action rewards minus a penalty for diverging from the prior policy.
@@ 1. AlphaZero’s MCTS-UCB
AlphaZero [silver2018general] is a method to train neural networks to play games at a high level. The main idea is to use Monte-Carlo tree search (MCTS) to explore the game tree and learn an improved value and policy for a given game state, where the exploration is based on prior values and policies from the current network. The network is trained on these improved values and policies, and becomes stronger over time.
@@ 2. Regularized Policy Optimization
An alternative to $\hat{\pi}$ was explored in “Monte-Carlo tree search as regularized policy optimization” [grill2020]. They define the optimized posterior policy $\bar{\pi}(s,\cdot)$ to be the one maximizing the expected reward (assuming estimated rewards $Q(s,a)$) minus a penalty for diverging from the prior policy $\pi_{0}(s,\cdot)$. Specifically, $\bar{\pi}(s,\cdot)$ is the unique distribution on actions at game state $s$ which maximizes
@@ 3. RMCTS
Recall that [grill2020] suggests replacing the posterior policy $\hat{\pi}$ with the optimized posterior policy $\bar{\pi}$, where the estimated $Q$ values are gotten from the original MCTS-UCB variant of AlphaZero. By contrast, RMCTS completely redefines MCTS itself, by using optimized policies not only at the root state, but throughout the tree. The $Q$-values are computed in a recursive manner, where the value of nodes $s^{\prime}$ in the search tree are approximated (recursively) by computing optimized posterior policies below $s^{\prime}$. See Algorithm 2 in section 4 for a succinct description of RMCTS.
@@ 8 $N(s,a_{i})\leftarrow\#\{k\in\mathbb{Z}:t_{i-1}\leq x+k<t_{i}\}$;
Output: Updated policy $\bar{\pi}$ as defined in [grill2020], maximizing $\sum_{a}\bar{\pi}(a)Q(a)-\frac{C}{\sqrt{N}}\operatorname{KL-Div}(\pi_{0}\mid\mid\bar{\pi})$
@@ 8. What next?
We emphasize that RMCTS is not limited to the specific optimal posterior policy of [grill2020] (Algorithm 4). Other variants of optimized posterior policies can be considered, for example we can replace the Kullback-Leibler divergence with other divergence measures. It would be interesting to see how well these variants perform. For example, if we reverse the order of $\pi_{0}$ and $\bar{\pi}$ in the Kullback-Leibler divergence: i.e. $\bar{\pi}$ maximizes
@@
Throughout this paper we have only talked about AlphaZero, and not mentioned its successor MuZero [schrittwieser2020]. However, RMCTS also applies to MuZero, since the search procedure of MuZero is once again MCTS-UCB, excepting that it is done in the simulated latent space of the learned model.

### v2

@@ Abstract.
We introduce a recursive AlphaZero-style Monte–Carlo tree search algorithm, “RMCTS”. The advantage of RMCTS over AlphaZero’s MCTS-UCB [3] is speed. In RMCTS, the search tree is explored in a breadth-first manner, so that network inferences naturally occur in large batches. This significantly reduces the GPU latency cost. We find that RMCTS is often more than 40 times faster than MCTS-UCB when searching a single root state, and about 3 times faster when searching a large batch of root states.
The recursion in RMCTS is based on computing optimized posterior policies at each game state in the search tree, starting from the leaves and working back up to the root. Here we use the posterior policy explored in “Monte–Carlo tree search as regularized policy optimization” [1]. Their posterior policy is the unique policy which maximizes the expected reward given estimated action rewards minus a penalty for diverging from the prior policy.
@@ 1. AlphaZero’s MCTS-UCB
AlphaZero [3] is a method to train neural networks to play games at a high level. The main idea is to use Monte-Carlo tree search (MCTS) to explore the game tree and learn an improved value and policy for a given game state, where the exploration is based on prior values and policies from the current network. The network is trained on these improved values and policies, and becomes stronger over time.
@@ 2. Regularized Policy Optimization
An alternative to $\hat{\pi}$ was explored in “Monte-Carlo tree search as regularized policy optimization” [1]. They define the optimized posterior policy $\bar{\pi}(s,\cdot)$ to be the one maximizing the expected reward (assuming estimated rewards $Q(s,a)$) minus a penalty for diverging from the prior policy $\pi_{0}(s,\cdot)$. Specifically, $\bar{\pi}(s,\cdot)$ is the unique distribution on actions at game state $s$ which maximizes
@@ 3. RMCTS
Recall that [1] suggests replacing the posterior policy $\hat{\pi}$ with the optimized posterior policy $\bar{\pi}$, where the estimated $Q$ values are gotten from the original MCTS-UCB variant of AlphaZero. By contrast, RMCTS completely redefines MCTS itself, by using optimized policies not only at the root state, but throughout the tree. The $Q$-values are computed in a recursive manner, where the value of nodes $s^{\prime}$ in the search tree are approximated (recursively) by computing optimized posterior policies below $s^{\prime}$. See Algorithm 2 in section 4 for a succinct description of RMCTS.
@@ 8 $N(s,a_{i})\leftarrow\#\{k\in\mathbb{Z}:t_{i-1}\leq x+k<t_{i}\}$;
Output: Updated policy $\bar{\pi}$ as defined in [1], maximizing $\sum_{a}\bar{\pi}(a)Q(a)-\frac{C}{\sqrt{N}}\operatorname{KL-Div}(\pi_{0}\mid\mid\bar{\pi})$
@@ 8. What next?
We emphasize that RMCTS is not limited to the specific optimal posterior policy of [1] (Algorithm 4). Other variants of optimized posterior policies can be considered, for example we can replace the Kullback-Leibler divergence with other divergence measures. It would be interesting to see how well these variants perform. For example, if we reverse the order of $\pi_{0}$ and $\bar{\pi}$ in the Kullback-Leibler divergence: i.e. $\bar{\pi}$ maximizes
@@
Throughout this paper we have only talked about AlphaZero, and not mentioned its successor MuZero [2]. However, RMCTS also applies to MuZero, since the search procedure of MuZero is once again MCTS-UCB, excepting that it is done in the simulated latent space of the learned model.